6.66 Length of job tenure. Researchers at the Ter ry Colle ege of us. the ber of Sam Business at the University of Georgia sampled 344 ness students and asked them this question: O course of your lifetime, what is the maximum num years you expect to work for any one employer? The sa ple resulted in 19.1 years. Assume that the sample ot students was randomly selected from the 6,000 undergrad. uate students at the Terry College and that ơ 6 years. a. Describe the sampling distribution of F. b. If the mean for the 6,000 undergraduate students is ver 18.5 years, find P(r> 19.1) c. If the mean for the 6,000 undergraduate students is μ 19.5 years, find P(F > 19.1) d. If P(F > 19.1)-.5, what is μ? e. If P(r 19.1) 2, is u greater than or less than 19.1 years? Explain. 6.67 Susceptibility to hypnosis. The Computer-Assisted

I need help with all parts but please explain in detail part d and e. Thank you!
Answer #1


The mean of the distribution of sample means is obtained below 19.1b. Find P>19.1) Formula:-x-μ /2 Substitute μ = 18.5, σ = 6 and n 344 P(T>19.1)-PIT> 19.1-18.51 344 0.6 0.3235 - P(Z >1.85) =1-P(z 1.85) = 1-0.96995 003 (from the standard normal table) Thus, the value of P(>19.1) is 0.03.o) Find P(x 19.1) Substitute μ = 195, σ = 6 and n = 344 P(r > 19.1)-PIT> 19.1-19.5 T344 -0.4 0.3235 = P(Z >-1.24) -1-P(Zs-1.2d. If P(T> 19.1)= 0.5, what is μ? From the given information, we observe that the value ofAx- Thus, the value of μ-1911e. If P(T> 19.1)=0.2 is μ greater than or less than 19.1 years? From the given information, the value of 19.1 falls to the right of the distribution. So 19.1 is to the right ofA,1L, μ < 19.1. Moreover, μ is less than 19.1 years.